a penny has mass of about 3g.compute the energy that would be released if it's rest mass were entirely converted into the energy (2.7×10^14J)

1
Expert's answer
2021-07-19T09:49:42-0400

Since we know that the mass-energy equivalence is "1\, kg \leftrightarrow 8.988\times 10^{16}\, J", we find:


"E=m_0c^2=(3\,\cancel{g}(\frac{1\,\cancel{kg}}{1000\,\cancel{g}})(\frac{8.988\times10^{16}\, J}{1\,\cancel{kg}}))
\\\implies E=2.6964\times 10^{14}\,J"


In conclusion, the energy that would be released (if its rest mass

were entirely converted into energy) is E 2.7 X 1014 J.


Reference:

  • Young, H. D., & Freedman, R. A. (2015). University Physics with Modern Physics and Mastering Physics. Academic Imports Sweden AB.